# Conformal prediction for uncertainties in nucleon-nucleon scattering

### v1

@@ I Introduction
Efforts have emerged to address these uncertainties Dobaczewski et al. (2014); Ireland and Nazarewicz (2015); Carlsson et al. (2016); Drischler et al. (2020); Duguet et al. (2024); Armstrong et al. (2025). One common approach is to analyze the residual dependence for various cutoffs, which can provide insight into the size of omitted higher-order terms Epelbaum et al. (2015). Recently, a Bayesian framework has been introduced for fitting the parameters that encode the impact of short-distance physics in chiral EFTs, where these parameters are treated as naturally of $\mathcal{O}(1)$ Schindler and Phillips (2009). This approach incorporates prior information through Bayes’ theorem, the principle of maximum entropy, and employs marginalization over “nuisance” parameters to construct reliable probability distributions for the parameters of interest. The Bayesian approach to UQ also helps identify which parts of the calculations (experiment, many-body methods, EFT Hamiltonians) need improvement to reduce the overall uncertainty Furnstahl et al. (2015a). These methods yield degree-of-belief (DOB) intervals and can incorporate prior knowledge and model structure Wesolowski et al. (2019); Furnstahl et al. (2015b); Melendez et al. (2017). Similarly, Gaussian Process (GP) models have been used to treat the expansion coefficients as smooth, correlated functions across inputs such as energy. GPs offer a flexible way to capture correlations and produce credible intervals, while also enabling model diagnostics Melendez et al. (2019). Though these approaches are powerful, they are not without limitations. Bayesian methods often depend on the choice of priors and assumptions about model correctness. The resulting uncertainty intervals may be sensitive to these choices and may not always provide guaranteed coverage. Likewise, GP-based methods rely on assumptions of smoothness and functional form, which may not hold in all cases.
@@ II Theory
As noted, conformal prediction provides a statistically rigorous, distribution-free framework for uncertainty quantification. When using a machine learning or statistical model, the predictions we make are based on the patterns observed in the training data. However, due to factors such as noisy data, inaccurate models, or limited information, these predictions will always include some inherent uncertainty. One of the key strengths of CP is its model-agnostic nature: it can be applied to a wide range of learning algorithms, including linear regression, or more complex models like Gaussian processes.
@@ II.1 Quantile Regression
In Fig. 1, the quantile function identifies the smallest value $y$ such that a certain percentage of data lies below it. For instance, the median corresponds to the $\tau=0.5$ quantile, meaning half of the values fall below it. This approach works well when we study the distribution of a single variable. However, when the distribution of $Y$ depends on another variable $X$, we estimate conditional quantiles, which describe how the quantiles of $Y$ change with some other variable $X$. Quantile regression allows us to estimate these conditional quantiles by modeling the $\tau$-quantile of $Y$ as a linear function of $X$, denoted as $Q_{Y}(\tau\mid X)=X^{T}\beta_{\tau}$.
@@
To illustrate this procedure, we will first show how quantile regression differs from the simple case of least-squares regression. In least squares, the objective is to estimate the coefficients $\beta$ that minimize the total squared distance between the observed data points $y_{i}$ and the predicted values $x_{i}^{T}\beta$, where $x_{i}$ is the feature vector for the $i$-th observation, and $n$ is the total number of data points:
$\min_{\beta}\sum_{i=1}^{n}(y_{i}-x_{i}^{T}\beta)^{2},$ (3)
@@
and $u$ is the residual or difference between the observed $y_{i}$ and predicted values $x_{i}^{T}\beta$. The loss function says that when $\tau=0.5$, positive and negative residuals are weighted the same, which means that we want the same number of data points above and below the line. For $\tau>0.5$, positive residuals are weighted more, meaning there would be more data points under the line. And for $\tau=0.1$, data points above the line are penalized more compared to those located below the line.
@@
$y_{i}=2+3x_{i}+x_{i}\varepsilon,\hskip 11.38092pt$ (6)
where $\varepsilon$ is normally distributed with mean 2 and standard deviation 2, implying a variance of Var($\varepsilon$) = $\sigma^{2}$ = $4$. Since the noise term is scaled by $x_{i}$, the variance of $x_{i}\varepsilon$ becomes Var($x_{i}\varepsilon$) = $\sigma^{2}x_{i}^{2}$ = $4x_{i}^{2}$. This input-dependent noise, i.e., heteroscedasticity, results in larger variability in y as x increases.
@@
Figure 2 shows the quantile regression fits, which are obtained by solving the minimization of Eq. (4), for various values of $\tau\in\{0.05,0.25,0.5,0.75,0.95\}$. This plot shows that the median regression line ($\tau=0.5$) divides the data points into two halves, while the lower ($\tau=0.05$) and upper ($\tau=0.95$) quantile lines capture the spread of the distribution. For instance, the quantile line corresponding to $\tau=0.95$ shows that for higher values of $y$, the rate of increase with respect to $x$ is significantly greater than for lower values of $y$. This visualization emphasizes the strength of quantile regression in capturing the behavior of data across different quantiles. In the next section, we will show how to use these quantiles to construct CP intervals.
@@ II.2 Conformal-Prediction Intervals
To see the strength of the conformal-prediction approach, suppose we have a dataset $(X_{i},Y_{i})$ (for $i=1,2,3,...,n$) drawn from an unknown distribution, where $X_{i}$ is the feature vector and $Y_{i}$ is the response variable. The goal here is not just to provide a point estimate $f(X_{n+1})$, but to construct a prediction set $C(X_{n+1})$, such that
@@
which states that the probability a new $Y_{n+1}$ is within the prediction set, generated by $C(X_{n+1})$, is greater than or equal to $1-\alpha$. This statement is known as the marginal coverage property Angelopoulos et al. (2025). The story regarding conditional inference guarantees is more complicated Barber et al. (2023). Here $\alpha\in(0,1)$ is a user-defined error level, and $1-\alpha$ is the desired coverage level. For example, if the user specifies $\alpha=0.05$, the method constructs prediction sets that contain the true $Y_{n+1}$ at least 95% of the time.
@@
We start by considering a simple, and somewhat naive, implementation of conformal prediction known as split conformal prediction Angelopoulos et al. (2025). In this method, data are divided into two subsets:
  1. 1.
    The training set is used to fit a predictive model $f$.
  2. 2.
    The calibration set is used to evaluate the model’s performance via conformity scores.
Conformity scores measure how well a predicted value aligns with the true observed value. A common conformity score is the absolute residual between the prediction and the observed data:
@@
    Determine the quantile of conformity scores $Q_{S}$($1-\alpha$) (meaning the ($1-\alpha$)-th quantile of conformity score $S$, which is different from that for the random variable $Y$).
@@
where we define:
$q=Q_{S}(1-\alpha)$ (10)
for simplicity of notation. If the data are independent and identically distributed, this method guarantees that the prediction interval will contain the true data $Y_{n+1}$, with probability $1-\alpha$, regardless of the underlying distribution.
@@
Figure 3 illustrates the naive 95% conformal-prediction interval constructed using the split conformal-prediction method applied to a toy dataset with heteroscedastic noise. Here, a simple linear model is trained on the training data, and the prediction interval is constructed based on the absolute residuals using a calibration set. As we can see, the resulting prediction band is symmetric around the model prediction and has a constant width across all $x$ values. This is because our choice for the conformity score, $S=|Y-{f(X)}|$, does not account for variations in the noise. While the absolute residual conformity score provides a simple and intuitive way to construct conformal prediction intervals, it always produces symmetric intervals, which may not be ideal when the data distribution is skewed or the variance is not constant. To address this, a more flexible approach is to use Conformalized Quantile Regression (CQR) which allows for asymmetric intervals Angelopoulos et al. (2025).
@@
which measures how far the true value $Y$ lies outside the predicted quantile interval at level $\alpha$. Next, we compute the $1-\alpha$ quantile of the conformity scores. Finally we define the CQR prediction interval for a new $X_{n+1}$ as:
@@
This conformal-prediction method adjusts the initial quantile interval by expanding it when $q$ is positive, or shrinking when $q$ is negative Angelopoulos et al. (2025). Figure 4 shows the 95% prediction band created using the CQR method. Compared to the standard residual-based conformal prediction, CQR produces asymmetric bands that adjust to the variability of the data. We can see that the band becomes wider as $x$ increases, which reflects the increasing noise.
Suppose we are given a set of posterior samples for a physical observable, such as a cross section, generated at a specific energy. Our goal is to construct a prediction interval that captures the true value with a desired coverage. To simulate this realistic setup, we use toy data sets drawn from three distinct distributions such as Gaussian, Student-t, and Chi-squared, then apply the conformal procedure to each. The data are randomly split into training, calibration, and test sets. The prediction model is defined as the mean of the training data:
@@
To build naive predicttion intervals using absolute residual conformity scores, we follow the steps outlined above. Then the naive prediction intervals for any new point can be calculated as:
$C=[\bar{Y}-q,\bar{Y}+q]$ (14)
In our applications, this procedure will be repeated independently for each energy level using the posterior samples.
Figure 5 compares the performance of the naive CP method under these distinct data distributions. For each case, the histogram represents the sample values, and the red line shows the true probability distribution function of each. The blue vertical dashed lines show the constructed 95% prediction interval. Despite the difference in the shape of the distributions, specially the skewness in the Chi-squared case, the conformal intervals still provide reliable coverage. This highlights one of CP’s most important features: it makes no assumptions about the underlying distribution, and yet provides uncertainty bands. The only problem is that, when we apply the naive CP to Chi-squared samples, the resulting prediction intervals can sometimes extend into negative values, which are not reasonable. To address this issue, we apply CQR, which adapts the prediction intervals to better respect the distribution’s shape. As shown in Fig. 6, using CQR ensures that the prediction intervals for the Chi-squared case remain non-negative, providing a meaningful uncertainty bands.
@@ III Applications
To quantify uncertainties in nucleon-nucleon observables such as the total cross section, we apply conformal prediction to construct uncertainty bands with guaranteed coverage. In order to test our approach, we start by considering posterior distributions of the total cross section, obtained either from the pointwise model or the Gaussian-Process (GP) model due to BUQEYE collaboration Melendez et al. (2019). We apply the CP approach to these posteriors to obtain distribution-free prediction intervals.
We then compute nucleon-nucleon phase shifts for realistic local nuclear Hamiltonians that are commonly used in many-body calculations such as Quantum Monte Carlo. We are able to use CP to generate uncertainty bands with guaranteed coverage across all partial waves and lab energies. We also discuss how we calculate the phase shifts as well as how our CP approach, which guarantees empirical coverage, could be used in the future when fitting these interactions.
@@ III.1 Pointwise Model
The pointwise model, as introduced in the BUQEYE collaboration Melendez et al. (2019), provides a novel method for quantifying truncation errors in chiral EFT observables. Here, we apply their ideas and terminology to generate posterior distributions of nucleon-nucleon cross sections. In chiral EFT, physical observables are expressed as an expansion in powers of $Q$, which is $y=y_{ref}\sum_{n=0}^{\infty}c_{n}Q^{n}.$
Figure 7: Posterior distributions of the expansion coefficient $c_{n}$ at the N2LO order, obtained from the pointwise model, for the total cross section at two energies (50 and 200 MeV), using the EKM potential for $R=0.9\ \text{fm}$. The distributions are created using 10000 samples.
@@
The posterior parameters $\nu$ and $\tau$ in the Student-t distribution are updated based on the known coefficients $\vec{c}_{k}$. They are computed as:
$\nu=\nu_{0}+n_{c},\hskip 28.45274pt\nu\tau^{2}=\nu_{0}\tau_{0}^{2}+\vec{c}_{k}^{~{}2}$ (16)
@@
where $\nu_{0}$ and $\tau_{0}$ are prior hyperparameters, and $n_{c}$ is the number of known coefficients in $\vec{c}_{k}$. Here, we use $\nu_{0}=1$ and $\tau_{0}=1$, and update $\nu$ and $\tau$ before sampling from the posterior.
@@
Figure 8: Posterior distributions for the total cross section at N2LO, obtained from the pointwise model using EKM potential for $R=0.9\ \text{fm}$ at energies 50 and 200 MeV. The histograms are created using 10000 posterior samples.
@@
Figure 9: CP intervals obtained for the total cross section $\sigma_{tot}$ at N2LO using the EKM potential at $R=0.9\ \text{fm}$, shown across a range of energies. The left panel corresponds to the pointwise model, while the right panel corresponds to the Gaussian-Process (GP) model. Dark and light blue bands denote the 68% and 95% CP intervals constructed from 10000 samples, respectively. In both cases, the width of the intervals increases with energy.
Figure 10: Comparison of 68% intervals for the total cross section $\sigma_{tot}$ at the N2LO order using the EKM potential at $R=0.9\ \text{fm}$ obtained from different statistical framework. The left panel shows results from the pointwise model, while the right panel corresponds to the Gaussian-Process model. In both panels, the blue shaded region represents the 68% CP interval, and the green hatched region denotes the 68% DOB interval as reported by BUQEYE in Ref. Melendez et al. (2017). The bands from the CP and BUQEYE procedures are in close agreement, indicating both approaches yield consistent uncertainty.
@@
To study the behavior of the expansion coefficient $c_{n}$, we draw samples from the Student-t posterior for each energy. Here, we take the data for nucleon-nucleon scattering, which are available on BUQEYE’s GitHub repository Scattering cross section data taken from . Figure 7 illustrates the resulting posterior distributions for $c_{n}$ at two energies (50 and 200 MeV), based on the EKM potential Epelbaum et al. (2015) with a cutoff of $R=0.9\ \text{fm}$. The blue histograms show the samples drawn from the pointwise posterior, and the red line represents the corresponding analytic Student-t probability density function. In terms of the EFT observables, the posterior distribution for the full prediction $y=y+y_{k}$ is Melendez et al. (2019),
$y\mid\vec{y}_{k},Q\sim t_{\nu}\left(y_{k},\ y_{\text{ref}}^{2}\frac{Q^{2(k+1)}}{1-Q^{2}}\tau^{2}\right)$ (17)
@@
Here, Q is again the expansion parameter and $y_{ref}$ is the reference scale. Next, we generate samples for the full observable $y$ at each energy using the posterior defined above. Figure 8 shows the distributions of the total cross section at the mentioned energies using EKM and for $R=0.9\ \text{fm}$. The histogram corresponds to the empirical density of the samples, and red curves indicate the analytic Student-t PDF. At 50 MeV, the distributions are narrow and roughly symmetric, centered around the predicted value $y_{k}$. This suggests that the uncertainty is small at low energies, where the expansion parameter Q is small. As we go to higher energies, especially at 200 MeV, the distribution is clearly broader. This indicates that the uncertainty becomes larger at higher energies.
@@
Figure 11: Histogram of conformity scores obtained from the pointwise model for the calibration set at two energies (50 and 200 MeV). The red dashed lines indicate the 95th percentile, and the black dashed line shows the median of the conformity scores.
@@
Figure 9(a) shows the CP intervals for the total cross section $\sigma_{tot}$ at N2LO order using the EKM potential with $R=0.9\ \text{fm}$ across a wide range of energies (we will discuss Fig. 9(b) in the following section). The dark and light blue bands represent the 68% and 95% CP intervals, respectively, and the black line represents the experimental results from the Nijmegen partial wave analysis Stoks et al. (1993). As the energy increases, the intervals become wider, reflecting the increasing uncertainty in the EFT predictions. At low energies, the cross sections are larger and the prediction intervals are relatively narrower. In contrast, at higher energies, the intervals broaden significantly, capturing the increased uncertainty. his shows how the CP method adaptively accounts for model uncertainty without relying on specific assumptions about the underlying distribution.
Figure 10 illustrates a side-by-side comparison of uncertainty bands for the total cross-section $\sigma_{tot}$ at N2LO, constructed using two different frameworks: conformal-prediction and BUQEYE’s Bayesian approach. The left and right panels show results from the pointwise and GP models, respectively. In each case, the 68% CP intervals are shown as blue shaded regions, while the BUQEYE 68% DOB intervals appear as green hatched bands. Notably, two methods yield remarkably similar intervals across the entire range of energies.
@@
Figure 11 shows how the conformity scores are distributed at different energies. The conformity score for each calibration point measures how far it is from the model prediction (based on the training set), and plays a key role in determining the final uncertainty interval. The black dashed line represents the median, and the red dashed line shows the 95th percentile of the conformity scores, denoted as $q$, which defines the width of the conformal prediction interval. As we can see, from 50 MeV up to 200 MeV, the tail of the distribution becomes longer and the value of $q$ increases. This reflects greater uncertainty in the data at higher energies. As a result, the prediction intervals become wider at higher energies.
To evaluate the validity of the conformal prediction intervals, we plot the empirical coverage to the target coverage levels across a range of $1-\alpha$. Figure 12 shows the results of this evaluation for the total cross section at $E=50\ \text{MeV}$, based on 4000 independent trials.

### v2

@@ I Introduction
Efforts have emerged to address these uncertainties Dobaczewski et al. (2014); Ireland and Nazarewicz (2015); Carlsson et al. (2016); Drischler et al. (2020); Duguet et al. (2024); Armstrong et al. (2025). One common approach is to analyze the residual dependence for various cutoffs, which can provide insight into the size of omitted higher-order terms Epelbaum et al. (2015). Recently, a Bayesian framework has been introduced for fitting the parameters that encode the impact of short-distance physics in chiral EFT, where these parameters are treated as naturally of $\mathcal{O}(1)$ Schindler and Phillips (2009). This approach incorporates prior information through Bayes’ theorem, the principle of maximum entropy, and employs marginalization over “nuisance” parameters to construct reliable probability distributions for the parameters of interest. The Bayesian approach to UQ also helps identify which parts of the calculations (experiment, many-body methods, EFT Hamiltonians) need improvement to reduce the overall uncertainty Furnstahl et al. (2015a). These methods yield degree-of-belief (DoB) intervals and can incorporate prior knowledge and model structure Wesolowski et al. (2019); Furnstahl et al. (2015b); Melendez et al. (2017). Similarly, Gaussian Process (GP) models have been used to treat the expansion coefficients as smooth, correlated functions across inputs such as energy. GPs offer a flexible way to capture correlations and produce credible intervals, while also enabling model diagnostics Melendez et al. (2019). Though these approaches are powerful, they are not without limitations. Bayesian methods often depend on the choice of priors and assumptions about model correctness. The resulting uncertainty intervals may be sensitive to these choices and may not always provide guaranteed coverage. Likewise, GP-based methods rely on assumptions of smoothness and functional form, which may not hold in all cases.
@@ II Theory
As noted, conformal prediction provides a statistically rigorous, distribution-free framework for uncertainty quantification. When using a machine-learning or statistical model, the predictions made are based on the patterns observed in the training data. However, due to factors such as noisy data, inaccurate models, or limited information, these predictions will always include some inherent uncertainty. One of the key strengths of CP is its model-agnostic nature: it can be applied to a wide range of learning algorithms, including linear regression, or more complex models like Gaussian processes.
@@ II.1 Quantile Regression
In Fig. 1, the quantile function identifies the smallest value $y$ such that a certain percentage of data lies below it. For instance, the median corresponds to the $\tau=0.5$ quantile, meaning half of the values fall below it. This approach works well when we study the distribution of a single variable. However, when the distribution of $Y$ depends on another variable $X$, we estimate conditional quantiles, which describe how the quantiles of $Y$ change with some other variable $X$. Quantile regression allows us to estimate these conditional quantiles by modeling the $\tau$-quantile of $Y$ as a linear function of $X$, denoted as $Q_{Y}(\tau\mid X)=f(X)_{\tau}$.
@@
To illustrate this procedure, we will first show how quantile regression differs from the simple case of least-squares regression. In least squares, the objective is to estimate some coefficients that minimize the total squared distance between the observed data points and the predicted values:
$\min{}\sum_{i=1}^{n}(y_{i}-f(x_{i}))^{2},$ (3)
@@
and $u$ is the residual or difference between the observed values and predicted values. The loss function says that when $\tau=0.5$, positive and negative residuals are weighted the same, which means that we want the same number of data points above and below the line. For $\tau>0.5$, positive residuals are weighted more, meaning there would be more data points under the line. And for $\tau=0.1$, data points above the line are penalized more compared to those located below the line.
@@
$Y=2+3X+X\varepsilon,\hskip 11.38092pt$ (6)
where $\varepsilon$ is normally distributed with mean 2 and standard deviation 2, implying a variance of Var($\varepsilon$) = $\sigma^{2}$ = $4$. Since the noise term is scaled by $X$, the variance of $X\varepsilon$ becomes Var($X\varepsilon$) = $\sigma^{2}X^{2}$ = $4X^{2}$. This input-dependent noise, i.e., heteroscedasticity, results in larger variability in $Y$ as $X$ increases.
@@
We show the quantile regression fits in Fig. 2, which are obtained by solving the minimization of Eq. (4), for various values of $\tau\in\{0.05,0.25,0.5,0.75,0.95\}$. The median regression line ($\tau=0.5$) divides the data points into two halves, while the lower ($\tau=0.05$) and upper ($\tau=0.95$) quantile lines capture the spread of the distribution. For instance, the quantile line corresponding to $\tau=0.95$ shows that for higher values of $y$, the rate of increase with respect to $x$ is significantly greater than for lower values of $y$. This visualization emphasizes the strength of quantile regression in capturing the behavior of data across different quantiles. In the next section, we will show how to use these quantiles to construct CP intervals.
@@ II.2 Conformal-Prediction Intervals
To see the strength of the conformal-prediction approach, suppose we begin with two random variables, $X$ and $Y$, where $X$ represents the independent variable and $Y$ the dependent variable. Their possible values are denoted by $x_{i}$ and $y_{i}$ drawn from an unknown distribution, for i=1,2,,ni=1,2,...,n. The goal here is not just to provide a point estimate $f(x_{n+1})$, but to construct a prediction set $C(x_{n+1})$, such that
@@
which states that the probability a new $y_{n+1}$ is within the prediction set, generated by $C(x_{n+1})$, is greater than or equal to $1-\alpha$. This statement is known as the marginal coverage property Angelopoulos et al. (2025). The story regarding conditional inference guarantees is more complicated Barber et al. (2023). Here $\alpha\in(0,1)$ is a user-defined error level, and $1-\alpha$ is the desired coverage level. For example, if the user specifies $\alpha=0.05$, the method constructs prediction sets that contain the true $y_{n+1}$ at least 95% of the time. In the following, we review the formalism introduced by Angelopoulos et al. (2025).
@@
We start by considering a simple, and somewhat naive, implementation of conformal prediction known as split conformal prediction. In this method, data are divided into two subsets: the training set, with $n_{\text{train}}$ data points, is used to fit a predictive model, and the calibration set, with $n_{\text{calib}}$ data points, is used to evaluate the model’s performance via conformity scores. Conformity scores measure how well a predicted value aligns with the true observed value. A common conformity score is the absolute residual between the prediction and the observed data:
@@
    Arrange these scores in ascending order, and define $q$ as the $\left[(1-\alpha)(n_{\text{calib}}+1)\right]$-th element of this ordered sequence.
@@
If the data are independent and identically distributed, this method guarantees that the prediction interval will contain the true data $y_{n+1}$, with probability $1-\alpha$, regardless of the underlying distribution.
@@
In Fig. 3 we plot the naive 95% conformal-prediction interval constructed using the split conformal-prediction method applied to a toy dataset with heteroscedastic noise. Here, a simple linear model is trained on the training data, and the prediction interval is constructed based on the absolute residuals using a calibration set. As we can see, the resulting prediction band is symmetric around the model prediction and has a constant width across all $x$ values. This is because our choice for the conformity score, S=|Yf(X)|S=|Y-{f(X)}|, does not account for variations in the noise. While the absolute residual conformity score provides a simple and intuitive way to construct conformal prediction intervals, it always produces symmetric intervals, which may not be ideal when the data distribution is skewed or the variance is not constant. To address this, a more flexible approach is to use Conformalized Quantile Regression (CQR) which allows for asymmetric intervals.
@@
which measures how far the true value $Y$ lies outside the predicted quantile interval at level $\alpha$. After computing $q$, using step 3 of the split conformal prediction procedure outlined above, finally define the CQR prediction interval for a new $x_{n+1}$ as:
@@
Where the training data points $n_{\text{train}}$ are used to estimate the lower and upper quantiles. This conformal-prediction method adjusts the initial quantile interval by expanding it when $q$ is positive, or shrinking when $q$ is negative. The 95% prediction band created using the CQR method in shown in Fig. 4. Compared to the standard residual-based conformal prediction, CQR produces asymmetric bands that adjust to the variability of the data. We can see that the band becomes wider as $x$ increases, which reflects the increasing noise.
Suppose we are given a set of posterior samples for a physical observable generated at a specific energy. Our goal is to construct a prediction interval that captures the true value with a desired coverage. To simulate this realistic setup, we use toy data sets drawn from three distinct distributions such as Gaussian, Student-t, and Chi-squared, then apply the conformal procedure to each. The prediction model is defined as the mean of the training data points:
@@
To build naive prediction intervals using absolute residual conformity scores, we follow the steps outlined above. Then the naive prediction intervals for any new point can be calculated as:
$C=[\bar{y}-q,\bar{y}+q].$ (14)
In Fig. 5 we compare the performance of the naive CP method under these distinct data distributions. For each case, the histogram represents the sample values, and the red line shows the true probability distribution function of each. The black vertical dashed lines show the constructed 95% prediction interval. Despite the difference in the shape of the distributions, notably the skewness in the Chi-squared case, the conformal intervals still provide reliable coverage. This highlights one of CP’s most important features: it makes no assumptions about the underlying distribution, and yet provides robust uncertainty bands. The only problem is that, when we apply the naive CP to Chi-squared samples, the resulting prediction intervals can sometimes extend into negative values, which are not reasonable. To address this issue, we again turn to CQR, which adapts the prediction intervals to better respect the distribution’s shape. As shown in Fig. 6, using CQR ensures that the prediction intervals for the Chi-squared case remain non-negative, providing a meaningful uncertainty band. In the following applications, we repeat the CQR procedure at each energy value, using the posterior samples to construct CP intervals.
@@ III Applications
To quantify uncertainties in nucleon-nucleon observables such as the total cross section, we apply conformal prediction to construct uncertainty bands with guaranteed coverage. In order to test our approach, we start by considering posterior distributions of the total cross section, obtained either from the pointwise model or the Gaussian-Process (GP) model, both due to BUQEYE collaboration Melendez et al. (2019). We apply the CQR approach to these posteriors to obtain distribution-free prediction intervals.
We then compute nucleon-nucleon phase shifts for realistic local nuclear Hamiltonians that are commonly used in many-body calculations such as Quantum Monte Carlo. We are able to use CP to generate uncertainty bands with guaranteed coverage across all partial waves and lab energies. We also discuss how we calculate the phase shifts as well as how our CQR approach, which guarantees empirical coverage, could be used in the future when fitting these interactions.
@@ III.1 Pointwise Model
The pointwise model, as introduced by the BUQEYE collaboration Melendez et al. (2019), provides a novel method for quantifying truncation errors in chiral EFT observables. Here, we apply their ideas and terminology to generate posterior distributions of nucleon-nucleon cross sections. In chiral EFT, physical observables are expressed as an expansion in powers of $Q$, which is $y=y_{ref}\sum_{n=0}^{\infty}c_{n}Q^{n}.$
Figure 7: Posterior distributions of the expansion coefficient $c_{n}$ at the N2LO order, obtained from the pointwise model, for the total cross section at two energies (50 and 200 MeV), using the EKM potential for $R_{0}=0.9\ \text{fm}$. The distributions are created using 10000 samples.
@@
The posterior parameters $\nu$ and $\tau$ in the Student-t distribution are updated based on the known coefficients $\bm{{c}_{k}}$. They are computed as :
$\nu=\nu_{0}+n_{c},\hskip 28.45274pt\nu\tau^{2}=\nu_{0}\tau_{0}^{2}+\bm{{c}_{k}}^{~2}$ (16)
@@
where $\nu_{0}$ and $\tau_{0}$ are prior hyperparameters, and $n_{c}$ is the number of known coefficients in $\bm{{c}_{k}}$. Here, we use $\nu_{0}=1$ and $\tau_{0}=1$, and update $\nu$ and $\tau$ before sampling from the posterior Melendez et al. (2019).
@@
Figure 8: Posterior distributions for the total cross section at N2LO, obtained from the pointwise model using EKM potential for $R_{0}=0.9\ \text{fm}$ at energies 50 and 200 MeV. The histograms are created using 10000 posterior samples.
@@
Figure 9: CQR intervals obtained for the total cross section $\sigma_{\text{tot}}$ at N2LO using the EKM potential at $R_{0}=0.9\,\text{fm}$, shown across a range of energies. The left panel corresponds to the pointwise model, while the right panel corresponds to the Gaussian-Process (GP) model. Dark and light blue bands denote the 68% and 95% CQR intervals constructed from 10000 samples, respectively. In both cases, the width of the intervals increases with energy. As we can see, the bands are jagged because of fluctuations in the underlying samples; this effect is stronger for the pointwise model, where the samples are noisier, and weaker for the GP, where the correlations across energies smooth behavior.
Figure 10: Empirical coverage distribution for the total cross section $\sigma_{tot}$ at N2LO and $E=50$ MeV using the EKM potential at $R_{0}=0.9\,\text{fm}$ for three different values of ($n_{\text{split}}=100,1000,10000$). The left panel shows the empirical coverage distribution for CQR using 2000 samples calculated from the GP, with $n_{\text{calib}}=900$, $n_{\text{train}}=900$, and $n_{\text{valid}}=200$. For each $n_{\text{split}}$, a new CQR interval is constructed, and this process is repeated $n_{\text{trial}}=1000$ times (where each trial uses freshly drawn samples). As $n_{\text{split}}$ increases, the spread of empirical coverage narrows and converges toward the target coverage $0.9$. The right plot illustrates the empirical distribution for DoB intervals, which remain fixed across different $n_{\text{split}}$, as they are not based on the calibration and training data points.
@@
To study the behavior of the expansion coefficient $c_{n}$, we draw samples from the Student-t posterior for each energy. Here, we take the data for nucleon-nucleon scattering, which are available on BUQEYE’s GitHub repository Scattering cross section data taken from . In Fig. 7 we plot the resulting posterior distributions for $c_{n}$ at two energies (50 and 200 MeV), based on the EKM potential Epelbaum et al. (2015) with a cutoff of $R_{0}=0.9\ \text{fm}$. The blue histograms show the samples drawn from the pointwise posterior, and the red line represents the corresponding analytic Student-t probability density function. In terms of the EFT observables, the posterior distribution for the full prediction $y=y_{k}+\delta y_{k}$ is Melendez et al. (2019),
$y\mid\bm{{y}_{k}},Q\sim t_{\nu}\left(y_{k},\ y_{\text{ref}}^{2}\frac{Q^{2(k+1)}}{1-Q^{2}}\tau^{2}\right)$ (17)
@@
Here, Q is again the expansion parameter and $y_{ref}$ is the reference scale. Next, we generate samples for the full observable $y$ at each energy using the posterior defined above. In Fig. 8 we show the distributions of the total cross section at the mentioned energies using EKM and for $R_{0}=0.9\ \text{fm}$. The histogram corresponds to the empirical density of the samples, and red curves indicate the analytic Student-t PDF. At 50 MeV, the distributions are narrow and roughly symmetric, centered around the predicted value $y_{k}$. This suggests that the uncertainty is small at low energies, where the expansion parameter Q is small. As we go to higher energies, especially at 200 MeV, the distribution is clearly broader. This indicates that the uncertainty becomes larger at higher energies.
@@
Figure 11: Histogram of the absolute residual conformity scores obtained from the pointwise model for the calibration set at two energies (50 and 200 MeV). The red dashed lines indicate the 95th percentile, and the black dashed line shows the median of the conformity scores.
@@
In Fig. 9 we show the CQR intervals for the total cross section $\sigma_{tot}$ at N2LO order using the EKM potential with $R_{0}=0.9\ \text{fm}$ across a wide range of energies (we will discuss the right panel of Fig. 9 in the following section). The dark and light blue bands represent the 68% and 95% CQR intervals, respectively, and the black line represents the experimental results from the Nijmegen partial wave analysis Stoks et al. (1993). As the energy increases, the intervals become wider, reflecting the increasing uncertainty in the EFT predictions. At low energies, the cross sections are larger and the prediction intervals are relatively narrower. In contrast, at higher energies, the intervals broaden significantly, capturing the increased uncertainty. This shows how the CQR method adaptively accounts for model uncertainty without relying on specific assumptions about the underlying distribution.
In Fig. 10 we show a side-by-side comparison of empirical coverage distributions for the total cross-section $\sigma_{tot}$ at N2LO and $E=50$ MeV for $n_{\text{split}}=100,1000,10000$, where $n_{\text{split}}$ refers to the number of times the dataset is randomly split into training, calibration, and validation sets. In the left panel we show the empirical coverage distributions for CQR. A dataset of 2000 samples was generated using a GP and divided into $n_{\text{train}}=900$, $n_{\text{calib}}=900$, and $n_{\text{valid}}=200$. For each value of $n_{\text{split}}$, a new CQR interval was formed, and the entire procedure was repeated $n_{\text{trial}}=1000$ times. Here $n_{\text{trial}}$ denotes the number of independent trials, each carried out with freshly drawn data from the GP. The results show that increasing $n_{\text{split}}$ reduces the spread of the empirical coverage distribution, which becomes narrower and centered around the target coverage $0.9$. In the right panel of Fig. 10 we show the empirical coverage distribution of the DoB under the same setup. Unlike the CQR results in the left panel, the DoB interval does not narrow with increasing $n_{\text{split}}$, since the DoB interval is fixed and does not depend on calibration data points. As a result, the three distributions overlap rather than getting narrower. This result demonstrates the validity guarantee promised by conformal prediction: as the number of times we’ve randomly split the data ($n_{\text{split}}$) increases, the empirical coverage distribution narrows toward the target coverage.
@@
In Fig. 11 we explore how the absolute residual conformity scores are distributed at different energies. The black dashed line represents the median, and the red dashed line shows the 95th percentile of the conformity scores, denoted as $q$, which defines the width of the conformal prediction interval. As we can see, from 50 MeV up to 200 MeV, the tail of the distribution becomes longer and the value of $q$ increases. This reflects greater uncertainty in the data at higher energies. As a result, the prediction intervals become wider at higher energies.